solution A is 40% acid. Solution B is 80% acid. How much of each is needed to make 120 liters of a solution that is 50% acid?
Answer by robertb(5830)

0.40*x + 0.80*(120 - x) = 0.50*120
==> 0.4x + 96 - 0.8x = 60
==> -0.4x = -36
==> x = 90 liters of 40% acid, and 120 - x = 30 liters 80% acid.